# Sparsification Under Siege: Defending Against Poisoning Attacks in Communication-Efficient Federated Learning

### v1

@@ Abstract
Federated Learning (FL) enables collaborative model training across distributed clients while preserving data privacy, yet it faces significant challenges in communication efficiency and vulnerability to poisoning attacks. While sparsification techniques mitigate communication overhead by transmitting only critical model parameters, they inadvertently amplify security risks: adversarial clients can exploit sparse updates to evade detection and degrade model performance. Existing defense mechanisms, designed for standard FL communication scenarios, are ineffective in addressing these vulnerabilities within sparsified FL. To bridge this gap, we propose FLARE, a novel federated learning framework that integrates sparse index mask inspection and model update sign similarity analysis to detect and mitigate poisoning attacks in sparsified FL. Extensive experiments across multiple datasets and adversarial scenarios demonstrate that FLARE significantly outperforms existing defense strategies, effectively securing sparsified FL against poisoning attacks while maintaining communication efficiency.
@@ I INTRODUCTION
Federated Learning (FL)[1] has emerged as a transformative paradigm for decentralized machine learning, enabling multiple clients to collaboratively train a shared global model while retaining their raw data locally. By maintaining data localized, FL enhances privacy and minimizes the risk of data leakage, making it highly applicable to privacy-sensitive domains such as healthcare, finance, and distributed IoT networks [2, 3, 4]. Despite these advantages, the iterative nature of FL necessitates frequent transmission of model parameters between clients and the central server, leading to substantial communication overhead[5]. This limitation is especially pronounced in resource-constrained environments, where bandwidth limitations and network instability can severely impact the scalability and practicality of the system.
@@
In contrast to standard FL, where robust aggregation methods such as Median[10], Trimmed Mean[10], and Krum[11] have demonstrated effectiveness in mitigating poisoning attacks, sparsified FL introduces distinct challenges. The selective transmission of updates can obscure malicious modifications, enabling adversarial clients to evade detection while exerting significant influence over the global model. These vulnerabilities highlight an urgent need for the development of defense mechanisms specifically tailored for sparsified FL, ensuring that communication efficiency achieved through sparsification does not compromise the system’s robustness against adversarial threats.
@@
  • We systematically analyze the security vulnerabilities of sparsified FL under poisoning attacks and demonstrate that existing defense mechanisms, originally designed for standard FL, are ineffective in sparsified settings.
  • We propose FLARE, a novel FL framework tailored for achieving robust aggregation against poisoning attacks while maintaining sparsified communication efficiency, which integrates sparse index mask inspection and model update sign similarity analysis to detect and mitigate adversarial behaviors.
  • We conduct extensive empirical evaluations on multiple datasets and across diverse adversarial scenarios. The results demonstrate that FLARE outperforms existing defense mechanisms, effectively safeguarding sparsified FL against various poisoning attacks while maintaining communication efficiency.
@@ II-B Model Poisoning Attacks and Defense Strategies in FL
To mitigate the impact of poisoning attacks in FL, various defense strategies have been proposed. These strategies can be broadly categorized into robust aggregation rules and anomaly detection techniques. Robust aggregation methods, such as Median [10] and Trimmed Mean [10], aim to improve the resilience of the global model by mitigating the impact of malicious updates during the aggregation process. Median aggregation selects the component-wise median of all client updates, making it robust to outliers and adversarial manipulations. Trimmed Mean excludes a fraction of the largest and smallest update values before computing the mean, effectively mitigating the impact of extreme deviations introduced by adversaries. Anomaly detection techniques take a different approach by identifying and filtering out malicious clients before aggregation. In addition, Krum[11], a widely used anomaly detection-based method, selects a single client update that is closest to its neighbors in Euclidean space, assuming that the majority of clients are honest. By prioritizing updates that exhibit consistency with the majority, Krum effectively reduces the impact of outliers. However, these defense methods were primarily designed for standard FL and have not yet addressed attacker detection in sparse training scenarios.
II-C Sparsified Privacy-Preserving FL
Recent studies [12, 22] have explored sparsification techniques to enhance communication efficiency and address privacy concerns in FL. However, these methods primarily integrate sparsification into privacy-preserving FL without addressing adversarial threats like poisoning attacks. For example, FedPHE [12] combines homomorphic encryption with sparsification to reduce communication costs while protecting client privacy by preventing the server from accessing individual updates. Similarly, Ergun et al. [22] propose a method that uses sparsification with Secure Multi-Party Computation (SMPC) to ensure data privacy by sharing only aggregated updates with the server.
Other studies [13, 23] have investigated sparsification techniques such as sampling the global model to counter adversarial updates or focusing on peer-to-peer frameworks. However, these approaches are not directly applicable to standard communication-efficient FL frameworks. This highlights an urgent need for poisoning-aware secure aggregation methods specifically designed for communication-efficient FL systems.
@@ IV-A Overview of FLARE
Figure 1: An illustration of the FLARE training process, incorporating client-side top-$k$ sparsification and server-side robust and sparsified aggregation.
Existing sparsified communication efficiency approaches have been observed to obscure malicious modifications in poisoning attack scenarios, enabling adversarial clients to evade detection (demonstrated and discussed in Section V-C). Consequently, existing defense mechanisms, originally designed for standard FL, become ineffective when applied to communication-efficient FL frameworks. These findings underscore the urgent need for security-aware sparsification techniques to ensure both robustness and communication efficiency in FL systems. To address this challenge, the goal of the proposed FLARE framework is to enhance robustness against untargeted poisoning attacks in the context of sparsified, communication-efficient FL.
@@ 16      Set $W_{G}(P_{i})=w_{i}^{agg}$;
V Experimental Evaluations
V-A Implementation and Default Setting
@@
Our proposed FLARE [^1] The code for FLARE will be made publicly available on GitHub upon the final decision of this paper. is implemented using PFLlib [27], an open-source federated learning (FL) framework.
@@
V-B Datasets and Baselines
@@
V-C Poisoning Attacks in Sparsified FL
@@
V-D Model Performance and Defense Effectiveness of FLARE
@@
V-E Impact of Attacker Ratio in FLARE
@@
V-F Impact of Top-k Sparsification Ratio in FLARE
@@
VI Conclusion

### v3

@@ Abstract
Federated Learning (FL) facilitates collaborative model training across decentralized clients while preserving local data privacy. However, FL systems still face critical challenges, notably communication inefficiency and vulnerability to poisoning attacks. Sparsification techniques, which selectively communicate essential model parameters, effectively reduce communication overhead yet inadvertently introduce new security risks, as adversarial clients can exploit sparse updates to evade detection and significantly degrade model performance. Existing security mechanisms, primarily designed for standard dense-update FL scenarios, fail to adequately address the vulnerabilities unique to sparsified FL. To tackle these challenges, we propose FLARE, a novel framework explicitly designed for robust aggregation and defense in sparsified FL settings. FLARE combines sparse index mask inspection and model update sign similarity analysis, effectively identifying and mitigating malicious client updates. Extensive experiments conducted on multiple datasets and under various adversarial conditions demonstrate that FLARE significantly outperforms state-of-the-art defense strategies, providing stronger resilience against poisoning attacks while simultaneously preserving the communication efficiency advantages of sparsification.
@@ I Introduction
Federated Learning (FL)[1] has emerged as a transformative paradigm for decentralized machine learning, enabling multiple clients to collaboratively train a shared global model while retaining their raw data locally. Within the context of service computing, FL facilitates distributed intelligent services by enabling model training across multiple stakeholders while maintaining service consumers’ data privacy. Such capabilities enhance trustworthiness and scalability in privacy-sensitive service sectors, including personalized healthcare services, financial analytics, and large-scale distributed IoT environments [2, 3, 4]. Despite these advantages, the iterative nature of FL necessitates frequent transmission of model parameters between clients and the central server, leading to substantial communication overhead[5]. This limitation is especially pronounced in resource-constrained environments, where bandwidth limitations and network instability can severely impact the scalability and practicality of the system.
@@
In contrast to standard FL, where robust aggregation methods such as Median[10], Krum[11], and RFA[12] have demonstrated effectiveness in mitigating poisoning attacks, sparsified FL introduces distinct challenges. The selective transmission of updates can obscure malicious modifications, enabling adversarial clients to evade detection while exerting significant influence over the global model. These vulnerabilities highlight an urgent need for the development of defense mechanisms specifically tailored for sparsified FL, ensuring that communication efficiency achieved through sparsification does not compromise the system’s robustness against adversarial threats.
@@
  • Vulnerability Analysis for Sparsified FL under Poisoning Attacks: We systematically analyze the security vulnerabilities of sparsified FL under poisoning attacks and demonstrate that existing defense mechanisms, originally designed for standard FL, are ineffective in sparsified settings.
  • Robust and Efficient Defense Mechanism: We propose FLARE, a novel FL framework tailored for achieving robust aggregation against poisoning attacks while maintaining sparsified communication efficiency, which integrates sparse index mask inspection and model update sign similarity analysis to detect and mitigate adversarial behaviors.
  • Comprehensive Evaluations: We conduct extensive empirical evaluations on multiple datasets and across diverse adversarial scenarios. The results demonstrate that FLARE outperforms existing defense mechanisms, effectively safeguarding sparsified FL against various poisoning attacks while maintaining communication efficiency.
Organization. In Section II, we discuss related work, and in Section III, we introduce the formulation and security model. We propose the FLARE framework in Section IV. Experimental evaluations are presented in Section VI. We conclude the paper in Section VII.
@@ II-B Model Poisoning Attacks and Defense Strategies in FL
To mitigate the impact of poisoning attacks in FL, various defense strategies have been proposed. These strategies can be broadly categorized into robust aggregation[10, 12] rules and anomaly detection techniques[11, 23, 24, 25, 26, 27, 28]. Robust aggregation methods, such as Median [10] and Trimmed Mean [10], aim to improve the resilience of the global model by mitigating the impact of malicious updates during the aggregation process. Median aggregation selects the component-wise median of all client updates, making it robust to outliers and adversarial manipulations. Trimmed Mean excludes a fraction of the largest and smallest update values before computing the mean, effectively mitigating the impact of extreme deviations introduced by adversaries. Anomaly detection techniques take a different approach by identifying and filtering out malicious clients before aggregation. In addition, Krum[11], a widely used anomaly detection-based method, selects a single client update that is closest to its neighbors in Euclidean space, assuming that the majority of clients are honest. By prioritizing updates that exhibit consistency with the majority, Krum effectively reduces the impact of outliers. However, these defense methods were primarily designed for standard FL and have not yet addressed attacker detection in sparse training scenarios.
@@ III-B Security Model
Figure 1: An illustration of the FLARE training process, incorporating client-side top-$k$ sparsification and server-side robust and sparsified aggregation. It also highlights the two key components of FLARE: Sparse Index Mask Inspection and Model Update Sign Similarity Analysis. The red sections indicate attackers, who may carry out data poisoning attacks (such as label flipping attack) or model poisoning attacks (including IPM, Gaussian, and scaling attacks) to degrade the performance of the global model.
@@ IV-A Overview of FLARE
Existing sparsified communication efficiency approaches have been observed to obscure malicious modifications in poisoning attack scenarios, enabling adversarial clients to evade detection (demonstrated and discussed in Section VI-C). Consequently, existing defense mechanisms, originally designed for standard FL, become ineffective when applied to communication-efficient FL frameworks. These findings underscore the urgent need for security-aware sparsification techniques to ensure both robustness and communication efficiency in FL systems. To address this challenge, the goal of the proposed FLARE framework is to enhance robustness against untargeted poisoning attacks in the context of sparsified, communication-efficient FL.
@@ 16      Set $W_{G}(P_{i})=w_{i}^{agg}$;
V Mathematical Analysis Of The Attack Under Sparsified Aggregation
In this section, we provide a theoretical analysis of the effectiveness of poisoning attacks within the context of the sparsified aggregation framework detailed in Section IV-C. Our analysis reveals a unique vulnerability introduced by sparsification: the ability for attackers to concentrate their influence on specific subsets of model parameters, a threat not as pronounced in standard dense-update FL.
V-A Definition of Attack Effectiveness
To quantify the impact of a poisoning attack, we must first define its effectiveness. In an untargeted poisoning attack, the adversary’s goal is to degrade the overall performance of the global model. A natural way to measure this degradation is to assess the deviation of the aggregated global model from an ideal, un-poisoned model.
Let $C$ be the set of all participating clients, which is a union of the set of benign clients $C_{B}$ and the set of malicious clients $C_{A}$, such that $C=C_{B}\cup C_{A}$. Following the robust aggregation process in Algorithm 2, let $W_{G}$ denote the final global model aggregated from all clients in $C$. We define an ideal global model, $W^{\prime}_{G}$, as the model that would have been aggregated if only the benign clients from the set $C_{B}$ had participated.
The effectiveness of the attack, denoted by $\rho$, is defined as the squared Euclidean distance between the parameters of the poisoned global model $W_{G}$ and the ideal benign model $W^{\prime}_{G}$:
$\rho=\|W_{G}-W^{\prime}_{G}\|_{2}^{2}$ (6)
This metric captures the total deviation induced by the malicious clients across all model parameters. A larger value of $\rho$ indicates a more successful and damaging attack.
V-B Bounding Attack
To analyze the attack effectiveness, we first formalize the sparsified aggregation process from a mathematical perspective. As described in Section IV-C, the aggregation is performed on a pack-by-pack basis. Let’s consider a single parameter pack $p$. The set of clients contributing to this pack is $S_{p}=\{i\in C\mid M_{i}(p)=1\}$, where $M_{i}(p)$ is the sparse index mask for client $i$ at pack $p$. The size of this set is $N_{p}=|S_{p}|$.
This set of contributors can be split into benign and malicious subsets: $S_{p}^{B}=S_{p}\cap C_{B}$ and $S_{p}^{A}=S_{p}\cap C_{A}$. Let their sizes be $N_{p}^{B}=|S_{p}^{B}|$ and $N_{p}^{A}=|S_{p}^{A}|$, respectively, with $N_{p}=N_{p}^{B}+N_{p}^{A}$.
For simplicity, we model the aggregation for pack $p$ (lines 10-14 in Algorithm 2) as a weighted average over the contributing clients. The poisoned global model parameter for pack $p$ is:
$W_{G}(p)=\frac{1}{N_{p}}\left(\sum_{j\in S_{p}^{B}}W_{j}(p)+\sum_{a\in S_{p}^{A}}W_{a}(p)\right)$ (7)
where $W_{j}(p)$ and $W_{a}(p)$ are the parameters for pack $p$ from a benign client $j$ and a malicious client $a$, respectively. For this analysis, we assume uniform weights for simplicity, though the principle holds for the weighted scheme in Algorithm 2.
The ideal benign model for the same pack $p$ would be aggregated only over the benign contributors:
$W^{\prime}_{G}(p)=\frac{1}{N_{p}^{B}}\sum_{j\in S_{p}^{B}}W_{j}(p)$ (8)
The core of the attack lies in the malicious parameters $W_{a}(p)$. We can model a malicious update as a deviation from the ideal benign update. We assume the attacker’s update for pack $p$ is bounded in its deviation from the ideal benign average for that pack. Let $\epsilon$ be the expected magnitude of perturbation for any parameter pack an attacker chooses to poison. This is formulated as:
$\mathbb{E}[\|W_{a}(p)-W^{\prime}_{G}(p)\|_{2}]\leq\epsilon$ (9)
for any attacker $a\in S_{p}^{A}$. This value $\epsilon$ represents the potency of the malicious craft. A larger $\epsilon$ signifies a more aggressive attack.
V-C Proof of the Attack Effectiveness
With the definitions above, we can now derive an upper bound for the attack effectiveness $\rho$. We start by analyzing the deviation for a single pack $p$.
$\begin{split}W_{G}(p)-W^{\prime}_{G}(p)&=\frac{1}{N_{p}}\left(\sum_{j\in S_{p}^{B}}W_{j}(p)+\sum_{a\in S_{p}^{A}}W_{a}(p)\right)\\ &\qquad-W^{\prime}_{G}(p)\\ &=\frac{1}{N_{p}}\left(N_{p}^{B}W^{\prime}_{G}(p)+\sum_{a\in S_{p}^{A}}W_{a}(p)\right)\\ &\qquad-W^{\prime}_{G}(p)\\ &=\left(\frac{N_{p}^{B}}{N_{p}}-1\right)W^{\prime}_{G}(p)+\frac{1}{N_{p}}\sum_{a\in S_{p}^{A}}W_{a}(p)\\ &=\frac{-N_{p}^{A}}{N_{p}}W^{\prime}_{G}(p)+\frac{1}{N_{p}}\sum_{a\in S_{p}^{A}}W_{a}(p)\\ &=\frac{1}{N_{p}}\sum_{a\in S_{p}^{A}}(W_{a}(p)-W^{\prime}_{G}(p))\end{split}$ (10)
This equation elegantly captures the error for a single pack: it is the average deviation of the malicious contributors from the ideal benign model, scaled by the malicious-to-total contributor ratio for that pack.
Now, we bound the squared norm of this deviation using the triangle inequality and our perturbation bound from Equation 9:
$\begin{split}\|W_{G}(p)-W^{\prime}_{G}(p)\|_{2}^{2}&=\left\|\frac{1}{N_{p}}\sum_{a\in S_{p}^{A}}(W_{a}(p)-W^{\prime}_{G}(p))\right\|_{2}^{2}\\ &\leq\frac{1}{N_{p}^{2}}\left(\sum_{a\in S_{p}^{A}}\|W_{a}(p)-W^{\prime}_{G}(p)\|_{2}\right)^{2}\\ &\leq\frac{1}{N_{p}^{2}}(N_{p}^{A}\epsilon)^{2}=\left(\frac{N_{p}^{A}}{N_{p}}\right)^{2}\epsilon^{2}\end{split}$ (11)
The total attack effectiveness $\rho$ is the sum of these squared deviations over all parameter packs:
$\rho=\sum_{p}\|W_{G}(p)-W^{\prime}_{G}(p)\|_{2}^{2}\leq\sum_{p}\left(\frac{N_{p}^{A}}{N_{p}}\right)^{2}\epsilon^{2}$ (12)
Let us define the malicious contributor ratio for pack $p$ as $f_{p}=N_{p}^{A}/N_{p}$. This ratio represents the fraction of clients that contributed to pack $p$ who were malicious. The final bound on the attack effectiveness is:
$\rho\leq\epsilon^{2}\sum_{p}f_{p}^{2}$ (13)
Implications. This result highlights a critical vulnerability unique to communication-efficient FL using sparsification. In standard FL, the overall attacker ratio $|C_{A}|/|C|$ dilutes the attack’s impact across all model parameters relatively uniformly. However, Equation 13 shows that in sparsified FL, the attack’s impact is determined by the malicious contributor ratio $f_{p}$ at each parameter pack. Adversaries can exploit this by coordinating their sparse index masks ($M_{a}$) to all contribute to a small, critical subset of parameter packs. For these targeted packs, they can drive $f_{p}$ close to 1, thus exerting immense influence and causing significant model deviation, even if the overall attacker ratio $|C_{A}|/|C|$ is low. This demonstrates that sparsification, while beneficial for efficiency, introduces a new attack vector that must be addressed by specialized defense mechanisms like FLARE.
VI Evaluations
VI-A Implementation and Default Setting
@@
Our proposed FLARE is implemented using PFLlib [32], an open-source federated learning (FL) framework.
@@
We evaluate FLARE under two common data partitioning settings in FL: IID and Non-IID. In the IID setting, the dataset is evenly distributed among clients, with each client receiving an equal share of samples from all classes to ensure uniformity. Conversely, the Non-IID setting simulates statistical heterogeneity using a Dirichlet distribution-based method [33], where a concentration parameter $\alpha$ determines the level of data heterogeneity.
VI-B Datasets and Baselines
@@
  • LASA [24]: An approach that combines pre-aggregation sparsification with layer-wise adaptive aggregation. It first sparsifies updates to reduce the impact of malicious parameters, then uses a layer-wise adaptive filter that leverages both magnitude and direction metrics to select and aggregate benign layers.
  • FLGuardian [23]: A layer-space defense method that detects benign clients for each layer using a combination of pairwise cosine/Euclidean distances and a clustering algorithm. It then assigns a trust score to each client and aggregates updates from clients with the highest scores.
  • RFA [12]: A robust aggregation approach that uses the geometric median of client updates to produce the global model. The geometric median is computed efficiently using a Weiszfeld-type algorithm.
@@
VI-C Poisoning Attacks in Sparsified FL
@@
The limited effectiveness of existing defense methods in sparsified FL stems from their incompatibility with sparse update patterns. Robust aggregation rules such as Median, Trimmed Mean, and Multi-KRUM rely on full, aligned model updates across clients to compute reliable statistics. However, sparsification leads to inconsistent parameter subsets, breaking these assumptions and making such methods unreliable or even inapplicable. Meanwhile, anomaly detection techniques like FedSIGN and Robust-DPFL typically extract global patterns from dense updates, but sparse representations obscure critical behaviors and reduce detection granularity. Moreover, adversaries can exploit the sparsity itself—selectively manipulating sparse indices—to evade these defenses. These challenges highlight the need for sparsity-aware mechanisms specifically tailored for robust federated learning under communication constraints.
Figure 8: Comparison of defense effectiveness across different $\beta$ settings, evaluated on CIFAR-100 under LFA, GNA, IPM attack, and scaling attack.
Figure 9: Comparison of defense effectiveness across different $\gamma$ settings, evaluated on CIFAR-100 under LFA, GNA, IPM attack, and scaling attack.
VI-D Model Performance and Defense Effectiveness of FLARE
@@
VI-E Impact of Attacker Ratio in FLARE
@@
VI-F Impact of Top-k Sparsification Ratio in FLARE
@@
VI-G Ablation Study on Filtering Hyperparameters
To further evaluate the robustness and adaptability of the Poisoning Client Filtering mechanism in FLARE (Algorithm 1), we conduct ablation experiments on its two key hyperparameters: the filtering threshold $\beta$ and the clustering sensitivity $\gamma$, as illustrated in Figure 8 and Figure 9, respectively.
We first assess the impact of the filtering threshold $\beta$, which controls the cutoff value for sparsity mask similarity when filtering out potentially poisoned clients. We experiment with five values: $\beta\in\{0.2,0.4,0.6,0.8,1.0\}$. As shown in our results, FLARE maintains effective defense performance under all settings, indicating its robustness to $\beta$ variations. However, higher values of $\beta$ (e.g., 0.8 or 1.0) lead to over-filtering, where benign clients are mistakenly excluded from training, reducing model diversity and performance. On the other hand, lower values retain more benign clients while still effectively filtering out attackers. We adopt $\beta=0.6$ in our main experiments as it strikes a good balance between defense strength and benign client retention.
Next, we examine the effect of the clustering sensitivity parameter $\gamma$, which influences the neighborhood size for DBSCAN-based client clustering. We vary $\gamma$ from 0.1 to 0.5 in increments of 0.05. The results reveal that excessively high values (e.g., $\gamma>0.4$) cause the clustering step to fail in separating attackers from benign clients, leading to ineffective defense.
VII Conclusion